OpenStudy (anonymous):

Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30. If x represents the number of quarters in the jar, which equation can be used to find the number of quarters? @Alyx_da_Boss help

OpenStudy (bakonloverk):

28.30 - 25x = 150 i think but im not sure

OpenStudy (anonymous):

x+y=150 .25x+.05y=28.30 Solve the first for y, and then substitute into the second: y=150-x .25x-.05(150-x)=28.30 I'll let you do the rest yourself, but x will eventually be 104

OpenStudy (anonymous):

thanks